# Complementary polynomials in quantum signal processing

### v1

@@ 4.5 Proof of Theorem 4
$\big{\lVert}|P(z)|^{2}-|(1-\tfrac{\varepsilon}{4})P(z)|^{2}\big{\rVert}_{\infty,\mathbb{T}}\leq\big{\lVert}P(z)+(1-\tfrac{\varepsilon}{4})P(z)\big{\rVert}_{\infty,\mathbb{T}}\cdot\big{\lvert}P(z)-(1-\tfrac{\varepsilon}{4})P(z)\big{\rVert}_{\infty,\mathbb{T}}<\frac{\varepsilon}{2}.$ (4.65)
@@
$\lVert Q(z)-\tilde{Q}_{2,N}(z)\rVert_{\infty,\mathbb{T}}<\frac{\varepsilon}{5}$ (4.67)
@@
$\big{\lVert}|Q(z)|^{2}-|\tilde{Q}_{2,N}(z)|^{2}\big{\rVert}_{\infty,\mathbb{T}}\leq\lVert 2Q(z)-Q(z)+\tilde{Q}_{2,N}(z)\rVert_{\infty,\mathbb{T}}\cdot\lVert Q(z)-\tilde{Q}_{2,N}(z)\rVert_{\infty,\mathbb{T}}\leq\bigg{(}2+\frac{\varepsilon}{5}\bigg{)}\frac{\varepsilon}{5}<\frac{\varepsilon}{2}.$ (4.68)
@@
The scaling (4.16) can be obtained by inserting $\varepsilon\to\varepsilon/5(d+1)$ and $\delta\to\varepsilon/4$ into (4.11).
@@ 5 Numerical results
In practice, our new algorithm for computation of complementary polynomials is extremely fast and accurate. A reference implementation of Algorithm 1 in Python using the FFT from the PyTorch library [24] is provided in Figure 1. While our proofs in the previous section focus on $N$ even, we run the algorithm for both odd and even choices of $N$; our results carry over. All benchmarks are performed on an M2 Macbook Pro with 16 GB RAM; we have not used any GPU acceleration that PyTorch optionally provides for the FFT.
To facilitate comparison with prior work, we compute the same loss function (4.2) as used in the optimization based method [3]. Here, we just compute the loss function $\tilde{\Phi}$ of the algorithm’s output as a measure of how well it satisfies the complementarity condition, see Proposition 4.1.1 and Corollary 3.1. In [3], the coefficients of an approximate complementary polynomial are determined by an optimization procedure minimizing the loss function.
@@
Figure 2(a) demonstrates Algorithm 1 for random polynomials with $\delta=0.2$, i.e. $\lVert P(z)\rVert_{\infty,\mathbb{T}}\leq 0.8$. We consider random polynomials up to degree $d=10^{7}$ and can numerically confirm the asymptotic scalings $N=O\left(d\log(1/\varepsilon)\right)$ from Theorem 3, up to logarithmic terms in $d$. It appears that the bound on $N$ in Theorem 3 is not tight and lower $N$ are sufficient in practice. Already at $N=4d$, our algorithm achieves a far better loss function than the optimization based method [3]. The runtime of Algorithm 1 is shown in Figure 3. We observe a runtime $O(N\log N)$ as expected from the FFTs. Comparing to [3]’s algorithm running on a CPU, we achieve far better runtimes for the same value of the loss function.
@@
Beyond its numerical efficacy, Algorithm 1 provably and reproducibly targets the same canonical solution, where $Q$ has no roots in $\mathbb{\mathbb{D}}$. This is in contrast to the optimization based algorithm from [3], which can (and does) converge to very different solutions $Q(z)$ for the same target polynomial $P(z)$ on successive runs. There is also no guarantee that the optimization doesn’t find a local minimum.
@@
For polynomials with $\delta=0$, $Q_{0}(z)\neq 1$ because $1-|P(z)|^{2}$ has roots on $z\in\mathbb{T}$. Theorem 4 proves that in those cases, Algorithm 2 provides a solution, scaling as $N=O\!\left(\frac{d}{\varepsilon}\log\frac{d}{\varepsilon}\right)$ for desired error $\varepsilon$. Instead of running Algorithm 2 including initial downscaling of the polynomial, here we simply run Algorithm 1 despite lack of rigorous justification and show the results in Figure 2(b). We are running the same code of the reference implementation (Figure 1) used for $\delta=0.2$. Even though Algorithm 1 assumes $Q_{0}(z)=1$, which is not the case any more, it still seems to give good results, even at degree $d=10^{7}$. In fact, we can still achieve losses better than those of [3] at significantly lower runtimes. The computation of $\log(1-|P(\omega_{N}^{n})|^{2})$ in step 2 of Algorithm 1 does not cause any issues, despite $1-|P(z)|^{2}$ having roots on $z\in\mathbb{T}$, because, generically, these will not be located exactly at roots of unity $\omega_{N}^{n}$ (otherwise, the polynomial can be rotated $P(z)\to P(\mathrm{e}^{i\alpha}z)$ by a suitable phase $\mathrm{e}^{i\alpha}$). Empirically we find a scaling of the algorithm as $N\approx O(1/\sqrt[4]{\varepsilon^{\prime}})$ for a desired accuracy $\varepsilon^{\prime}$ in the loss function $\tilde{\Phi}$; proving this could be the subject of future work.
@@
(a) Polynomials with $\delta=0.2$. We observe the approximate scaling $N\approx O(d\log 1/\tilde{\Phi})$, linear scaling with the degree $d$ and logarithmic scaling with the accuracy $\tilde{\Phi}$. This is commensurate with our proof of error scaling (Theorem 3) for Algorithm 1, up to logarithmic factors in $d$.
@@
(b) Polynomials with $\delta=0$. Despite lack of rigorous justification for Algorithm 1 in this regime (necessitating the use of Algorithm 2, slightly scaling down the polynomial), we observe that Algorithm 1 is efficacious with a scaling behaviour of $N\approx O(d/\sqrt[4]{\tilde{\Phi}})$ in the accuracy $\tilde{\Phi}$.
@@
Figure 2: Finding the complementary polynomial with the reference implementation (Figure 1) of Algorithm 1 for random polynomials of various degrees $d$. We plot the achieved loss function (4.2) by chosen FFT dimension $N$. The optimization code from GQSP [3] only reaches losses up to $10^{-6}$, which Algorithm 1 achieves at far faster runtimes (see Figure 3). In the plots, loss functions saturate around $10^{-14}$ due to double-precision floating-point arithmetics.
@@
Figure 3: Runtime of the reference implementation (Figure 1) of Algorithm 1 on an M2 Macbook Pro with 16GB RAM. We observe the scaling $N\log N$ due to the fast Fourier transformations.
@@ 6 Discussion
In this paper, we have addressed the solvability of the complementary polynomials problem, Problem 1. Our main mathematical result, Theorem 2, is an exact representation, written as a set of contour integrals, for the complementary polynomial throughout the entire complex plane. We use a Fourier analytic variant of Theorem 2, Corollary 2.2, as a basis for developing a numerical method to obtain the complementary polynomial explicitly in the monomial basis.
@@
Acknowledgements
We thank Earl Campbell, György Geher, Tharon Holdsworth, Tanuj Khattar, Nicholas Rubin, and especially Hari Krovi for useful discussions. This work was partially funded by Innovate UK (grant reference 10071684).
Appendix A QSP conventions

### v2

@@ 4.5 Proof of Theorem 4
$\big{\lVert}|P(z)|^{2}-|(1-\tfrac{\varepsilon}{4})P(z)|^{2}\big{\rVert}_{\infty,\mathbb{T}}\leq\big{\lVert}P(z)+(1-\tfrac{\varepsilon}{4})P(z)\big{\rVert}_{\infty,\mathbb{T}}\big{\lVert}P(z)-(1-\tfrac{\varepsilon}{4})P(z)\big{\rVert}_{\infty,\mathbb{T}}<\frac{\varepsilon}{2}.$ (4.65)
@@
$\big{\lVert}Q(z)-\tilde{Q}_{2,N}(z)\big{\rVert}_{\infty,\mathbb{T}}<\frac{\varepsilon}{5}$ (4.67)
@@
$\big{\lVert}|Q(z)|^{2}-|\tilde{Q}_{2,N}(z)|^{2}\big{\rVert}_{\infty,\mathbb{T}}\leq\big{\lVert}2Q(z)-Q(z)+\tilde{Q}_{2,N}(z)\big{\rVert}_{\infty,\mathbb{T}}\big{\lVert}Q(z)-\tilde{Q}_{2,N}(z)\big{\rVert}_{\infty,\mathbb{T}}\leq\bigg{(}2+\frac{\varepsilon}{5}\bigg{)}\frac{\varepsilon}{5}<\frac{\varepsilon}{2}.$ (4.68)
@@
The scaling (4.16) can be obtained by making the replacements $\varepsilon\to\frac{\varepsilon}{5(d+1)}$ and $\delta\to\frac{\varepsilon}{4}$ in (4.11).
@@ 5 Numerical results
In practice, our new algorithm for computation of complementary polynomials is extremely fast and accurate. A reference implementation of Algorithm 1 in Python using the FFT from the PyTorch library [23] is provided in Figure 1. First, we run our reference implementation for random polynomials (Section 5.1) and compare with prior work [3] using random polynomials. Then, we turn to practically useful examples of polynomials occurring in quantum algorithms. These include Hamiltonian simulation (Section 5.2), eigenvalue filtering (Section 5.3), and the sign function (Section 5.4). Also in these cases, our algorithm works very well; the achievable degrees are only limited by our ability to compute the approximant $P$ to the desired target function. With random polynomials we can exhibit our algorithm for far higher degrees.
All benchmarks are performed on an M2 Macbook Pro with 16 GB RAM; we have not used any GPU acceleration that PyTorch optionally provides for the FFT. We compute the complementary polynomial using our algorithm for various dimensions $N$ of the FFT. To evaluate the accuracy of the output $\tilde{Q}_{2,N}$, we compute the loss function $\tilde{\Phi}(P,\tilde{Q}_{2,N})$ (4.2), which is a measure of how well the algorithm output satisfies the complementarity condition (1.1); see Proposition 4.1.1 and Corollary 3.1.
For the random polynomials, we work in single-precision (complex64 data type) arithmetic to facilitate comparison with the method and results presented in [3]. In the practical examples we study, we work in double-precision (complex128 data type) arithmetic, demonstrating that Algorithm 1 can achieve errors as low as $10^{-30}$.
5.1 Random Polynomials
Here, we generate random polynomials $P$. The real and imaginary parts of each coefficient are independently sampled from a normal distribution of unit variance. Subsequently, we scale each polynomial to achieve $\lVert P(z)\rVert_{\infty,\mathbb{T}}=1-\delta$ for choices $\delta=0.2$ and $\delta=0$. While our proofs in the previous section focus on $N$ even, we run the algorithm for both odd and even choices of $N$; our results carry over.
@@ 5.1 Random Polynomials
Figure 2(a) exhibits Algorithm 1 on random polynomials with $\delta=0.2$, i.e., $\lVert P(z)\rVert_{\infty,\mathbb{T}}\leq 0.8$. We consider random polynomials up to degree $d=10^{7}$ and can numerically confirm the scaling $N=O\left(d\log\frac{1}{\varepsilon}\right)$ from Theorem 3, up to logarithmic terms in $d$. It appears that the bound on $N$ in Theorem 3 is not tight and lower $N$ are sufficient in practice. For practical applications we therefore suggest increasing $N$ until the desired accuracy is reached, rather than using 4.8.
The optimization-based method [3] uses the same loss function $\tilde{\Phi}$ (4.1). The best loss it achieves is indicated by a horizontal dashed line in Figure 2. Already at $N=4d$, our algorithm achieves a far better loss function. The runtime of Algorithm 1 is shown in Figure 3. We observe a runtime $O(N\log N)$, as expected from an FFT-based algorithm. In comparison with the algorithm of [3] run on a CPU, we achieve far better runtimes for the same values of the loss function.
@@
Beyond its numerical efficacy, Algorithm 1 provably and reproducibly targets the same canonical solution, where $Q$ has no roots in $\mathbb{\mathbb{D}}$. This is in contrast to the optimization based algorithm of [3], where the polynomial coefficients are optimized with respect to the loss function; this optimization can (and does) converge to different solutions $Q$ for the same target polynomial $P$ on successive runs. There is also no guarantee that the optimization does not converge to a local minimum.
@@
For polynomials with $\delta=0$, $Q_{0}(z)\neq 1$ because $1-|P(z)|^{2}$ has roots on $z\in\mathbb{T}$. Theorem 4 proves that in those cases, Algorithm 2 provides a solution, scaling as $N=O\!\left(\frac{d}{\varepsilon}\log\frac{d}{\varepsilon}\right)$ for desired error $\varepsilon$. Instead of running Algorithm 2, including initial downscaling of the polynomial, here we simply run Algorithm 1 despite lack of rigorous justification and show the results in Figure 2(b). We are running the same code of the reference implementation (Figure 1) used for $\delta=0.2$. Even though Algorithm 1 assumes $Q_{0}(z)=1$, which is not the case any more, it still seems to give good results, even at degree $d=10^{7}$. In fact, we can still achieve losses better than those of [3] at significantly lower runtimes. The computation of $\log(1-|P(\omega_{N}^{n})|^{2})$ in step 2 of Algorithm 1 does not cause any issues, despite $1-|P(z)|^{2}$ having roots on $z\in\mathbb{T}$, because, generically, these will not be located exactly at roots of unity $\omega_{N}^{n}$ (otherwise, the polynomial can be rotated $P(z)\to P(\mathrm{e}^{\mathrm{i}\alpha}z)$ by a suitable phase $\mathrm{e}^{\mathrm{i}\alpha}$). Empirically, by performing a fit on the data in the figure, we find a scaling of the algorithm as $N=O\big{(}\frac{1}{\sqrt[4]{\varepsilon}}\big{)}$ for a desired accuracy $\tilde{\varepsilon}$ in the loss function $\tilde{\Phi}$; proving this scaling could be the subject of future work.
@@
(a) Polynomials with $\delta=0.2$. We observe the scaling $N=O\big{(}d\log\frac{1}{\tilde{\Phi}}\big{)}$, linear scaling with the degree $d$ and logarithmic scaling with the accuracy $\tilde{\Phi}$. This is commensurate with our proof of error scaling (Theorem 3) for Algorithm 1, up to logarithmic factors in $d$.
@@
(b) Polynomials with $\delta=0$. Despite lack of rigorous justification for Algorithm 1 in this regime (necessitating the use of Algorithm 2, slightly scaling down the polynomial), we observe that Algorithm 1 is efficacious with a scaling behavior of $N=O\big{(}\frac{d}{\sqrt[4]{\tilde{\Phi}}}\big{)}$ in the accuracy $\tilde{\Phi}$.
@@
Figure 2: Finding the complementary polynomial with the reference implementation (Figure 1) of Algorithm 1 for random polynomials of various degrees $d$; see Section 5.1. We plot the achieved loss function (4.2) by chosen FFT dimension $N$. The optimization code from GQSP [3] only reaches losses up to $10^{-6}$, which Algorithm 1 achieves at far faster runtimes; see Figure 3. In the plots, the loss functions saturate around $10^{-14}$ due to single-precision floating-point arithmetic.
@@
Figure 3: Runtime of the reference implementation (Figure 1) of Algorithm 1 on an M2 Macbook Pro with 16GB RAM for random polynomials; see Section 5.1. We observe the scaling $N\log N$ due to the FFTs.
5.2 Hamiltonian simulation
The Jacobi-Anger expansion reads
$\mathrm{e}^{-\mathrm{i}\tau x}=J_{0}(\tau)+2\sum_{n=1}^{\infty}\mathrm{i}^{n}J_{n}(\tau)T_{n}(x)\quad(\tau\in[0,\infty),x\in[-1,1])$ (5.1)
where $\{J_{n}(\tau)\}_{n=0}^{\infty}$ are Bessel functions of the first kind [24, Chapter 10] and $\{T_{n}(x)\}_{n=0}^{\infty}$ are Chebyshev polynomials of the first kind [24, Chapter 18]. In quantum algorithms, the QSVT is used to apply this function to a Hamiltonian to construct the time evolution operator [4, 6].
Denote the truncation of the Chebsyhev series (5.1) by
$f_{M}(x;\tau)\coloneqq J_{0}(\tau)+2\sum_{n=1}^{M}\mathrm{i}^{n}J_{n}(\tau)T_{n}(x)\quad(x\in[-1,1]).$ (5.2)
It may be shown that [18]
$\big{\lVert}f_{M}(x;\tau)-\mathrm{e}^{-\mathrm{i}\tau x}\big{\rVert}_{\infty,[-,1,1]}<\mathrm{e}^{\frac{1}{2}\mathrm{e}\tau-M}.$ (5.3)
Thus, choosing $M=\big{\lceil}\tfrac{1}{2}\mathrm{e}\tau+\log\tfrac{1}{\varepsilon}\big{\rceil}$, we are guaranteed that the polynomial
$\tilde{f}_{M}(x;\tau)\coloneqq\frac{1}{1+\varepsilon}f_{M}(x;\tau)$ (5.4)
$\varepsilon$-approximates $\mathrm{e}^{-\mathrm{i}\tau x}$ on $[-1,1]$ and satisfies $\lVert\tilde{f}_{M}(\tau;z)\rVert_{\infty,[-1,1]}<1$. The corresponding polynomial on $\mathbb{T}$ is given by $P(z)=z^{M}\tilde{f}_{M}(\tfrac{1}{2}(z+z^{-1});\tau)$; see Appendix A. We show results of Algorithm 1 in Figure 4, which can accurately compute the complementary polynomial with only very low overhead in FFT dimension $N$.
Figure 4: Computing the complementary polynomial with Algorithm 1 for Hamiltonian Simulation polynomials; see Section 5.2.
5.3 Eigenvalue filtering
The polynomial defined by
$g_{M}(x;a)\coloneqq\frac{T_{M}(\frac{2x^{2}-(1+a^{2})}{1-a^{2}})}{T_{M}(-\frac{1+a^{2}}{1-a^{2}})}\quad(a\in(0,1)).$ (5.5)
is used in quantum algorithms for eigenvalue filtering [25]. It has a sharp peak at $x=0$; see Figure 5(a) for an example. In quantum algorithms it can be used to project onto the kernel of a matrix.
(a) Example polynomial for eigenvalue filtering [25].
(b) Computing the complementary polynomial with Algorithm 1.
Figure 5: Eigenvalue filtering polynomial; see Section 5.3.
While the expansion of (5.5) in the Chebyshev basis can be found explicitly via standard identities, it leads to expressions for coefficients that are numerically unstable. Instead, the Chebsyhev coefficients of (5.5) may be determined via a Chebyshev transform, as we now describe. Setting
$x_{m,M}\coloneqq\cos\bigg{(}\frac{2m+1}{2M}\pi\bigg{)},$ (5.6)
and using the discrete orthogonality of the Chebsyhev polynomials, it may be shown that
$g_{M}(x;a)=c_{0}+\sum_{m=1}^{2M}c_{m}T_{m}(x),$ (5.7)
where
$c_{0}=\frac{1}{2M+1}\sum_{m=0}^{2M}g_{M}(x_{m,2M+1};a),\quad c_{n}=\frac{2}{2M+1}\sum_{m=0}^{2M}g_{M}(x_{m,2M+1};a)T_{m}(x_{m,2M+1}).$ (5.8)
We perform the Chebyshev transformation with the Chebyshev.interpolate() function in NumPy [26]. The corresponding polynomial on $\mathbb{T}$ is $P(z)=z^{M}g_{M}\big{(}\tfrac{1}{2}\big{(}z^{\frac{1}{2}}+z^{-\frac{1}{2}}\big{)};a\big{)}$; see Appendix A.
In order to run our reference implementation of Algorithm 1, we multiply the polynomial $R_{M}(z;a)$ by $(1-10^{-10})$. The results in Figure 5(b) demonstrate that our algorithm works very well.
5.4 Signum function
Uniform polynomial approximations of the signum function, defined by
$\mathrm{sgn}(x)\coloneqq\begin{cases}-1&x<0\\ 0&x=0\\ +1&x>0,\end{cases}$ (5.9)
are required in various QSVT-based applications including amplitude amplification and phase estimation [6]. As $\mathrm{sgn}(x)$ is not regular at $x=0$, a standard approach to the construction of polynomial approximants uses an error function with an appropriately rescaled argument as a regularization of the signum function. The resulting polynomial approximant is [27]
$h_{M}(x;\beta)\coloneqq 2\mathrm{e}^{-\beta}\sqrt{\frac{2\beta}{\pi}}\Bigg{(}I_{0}(\beta)T_{1}(x)+\sum_{n=1}^{M}(-1)^{n}I_{n}(\beta)\bigg{(}\frac{T_{2n+1}(x)}{2n+1}-\frac{T_{2n-1}(x)}{2n-1}\bigg{)}\Bigg{)},$ (5.10)
where $\{I_{n}(\beta)\}_{n=0}^{\infty}$ are modified Bessel functions [24, Chapter 10].
The following is a essentially a variant of [27, Theorem 3], convenient for our purposes. Below, $W_{0}$ denotes the principal branch of the Lambert $W$-function [24, Section 4.13].
Theorem 5 (Chebyshev approximation of the signum function, [27]).
For any $a\in(0,1)$ and $\varepsilon\in\big{(}0,\frac{3}{\sqrt{8\pi\log 2}}\big{)}$, let $\beta>0$ satisfy $\beta\geq\frac{1}{4a^{2}}W_{0}\big{(}\frac{18}{\pi\varepsilon^{2}}\big{)}$ and $M\in\mathbb{Z}_{\geq 1}$ satisfy
$M\geq\sqrt{\frac{W_{0}\big{(}\frac{72}{\pi\varepsilon^{2}}\big{)}\bigg{(}\log\bigg{(}\frac{3}{\sqrt{2\pi}}\frac{1}{\varepsilon\sqrt{W_{0}\big{(}\frac{72}{\pi\varepsilon^{2}}\big{)}}}\bigg{)}-\beta\bigg{)}}{W_{0}\bigg{(}\frac{1}{\mathrm{e}}\bigg{(}\frac{1}{\beta}\log\bigg{(}\frac{3}{\sqrt{2\pi}}\frac{1}{\varepsilon\sqrt{W_{0}\big{(}\frac{72}{\pi\varepsilon^{2}}\big{)}}}\bigg{)}-1\bigg{)}\bigg{)}}}.$ (5.11)
Then,
$\lVert\mathrm{sgn}(x)-h_{M}(x;\beta)\rVert_{\infty,[-1,-a]\cup[a,1]}<\varepsilon$ (5.12)
and
$\lVert h_{M}(x;\beta)\rVert_{\infty,[-1,1]}<1+\tfrac{2}{3}\varepsilon.$ (5.13)
hold.
Based on Theorem 5, we consider the polynomial
$\tilde{h}_{M}(x;\beta)\coloneqq\frac{1}{1+\tfrac{2}{3}\varepsilon}h_{M}(x;\beta)$ (5.14)
which $\varepsilon$-approximates the signum function on $[-1,-a]\cup[a,1]$ and is strictly bounded in absolute value by unity on $[-1,1]$. The corresponding polynomial on $\mathbb{T}$ is $P(z)=z^{(2M+1)/2}\tilde{h}_{M}\big{(}\tfrac{1}{2}\big{(}z^{\tfrac{1}{2}}+z^{-\tfrac{1}{2}}\big{)};\beta\big{)}$; see Appendix A.
To test our algorithm, we use example polynomials given by the parameters in Figure 6(a). Our results in Figure 6(a) show that Algorithm 1 works very well in practice with only small overhead in the FFT dimension $N$.
Example Specification of polynomial
$a$ $\varepsilon$ $\beta$ $M$
1 $10^{-1}$ $10^{-1}$ 120 29
2 $10^{-1}$ $10^{-4}$ 433 99
3 $10^{-1}$ $10^{-7}$ 765 172
4 $10^{-1}$ $10^{-10}$ 1101 246
5 $10^{-4}$ $10^{-1}$ 119631742 26690
6 $10^{-4}$ $10^{-4}$ 432869078 89806
7 $10^{-4}$ $10^{-7}$ 764051835 156148
(a) Parameters of example signum function polynomials; see Theorem 5.
(b) Computing the complementary polynomial with Algorithm 1.
Figure 6: Signum function polynomial; see Section 5.4.
@@ 6 Discussion
In this paper, we have addressed the analytic and numerical solvability of the complementary polynomials problem, Problem 1. Our main mathematical result, Theorem 2, is an exact representation, written as a set of contour integrals, for the complementary polynomial throughout the entire complex plane. We use a Fourier analytic variant of Theorem 2, Corollary 2.2, as a basis for developing a numerical method to obtain the complementary polynomial explicitly in the monomial basis.
@@
  7. 7.
    After this paper was posted to the arXiv, several related works have appeared. In [33], an FFT-based algorithm for computing individual phase factors is developed. More specifically, [33] presents a similar method to Algorithm 1 for computing complementary polynomials as a subroutine in their algorithm, and subsequently shows how to compute individual phase factors via an FFT and linear algebra. Improvements to the linear algebraic component of the algorithm were reported in [34]. In [35, 36], a Newton-Raphson-based algorithm for computing complementary polynomials is introduced and analyzed. Numerical experiments performed on our algorithm in [35] overlap with the examples in Sections 5.15.2.
  8. 8.
    Qualtran [37] is a recently-introduced platform for quantum algorithm development. As described in [37], Algorithm 2 has been integrated into Qualtran for the purpose of constructing QSP and QSVT circuits.
Acknowledgements.
We thank Earl Campbell, György Geher, Tharon Holdsworth, Tanuj Khattar, Edwin Langmann, Nicholas Rubin, and especially Hari Krovi for useful discussions. We are grateful to Nina Glaser for bringing our attention to [25] and suggesting eigenvalue filtering as an example. We thank three anonymous referees for suggestions that improved the paper. This work was partially funded by Innovate UK (grant reference 10071684).
Data Availability.
Data sets generated during the course of this study will be made available upon reasonable request.
Conflict of interest.
On behalf of all authors, the corresponding author states that there is no conflict of interest.
Appendix A Quantum signal processing conventions